OpenStudy (anonymous):

Which equation would best help solve the following problem: Gary kicks a field goal with an initial vertical velocity of 38 meters per second. How long will it take the football to hit the ground?

OpenStudy (anonymous):

9t2 + 42t = 0

OpenStudy (anonymous):

let me give you the choices

OpenStudy (anonymous):

–16t2+ 38t = 0 16t2 + 38t = 0 –4.9t2 + 38t = 0 –4.9t2 – 38t = 0

OpenStudy (anonymous):

it's c